OpenStudy (anonymous):

find the indefinite integral using substitutions: x sin 3x^2 dx

OpenStudy (anonymous):

let u=3x^2, du = 6x so...our integral is \[1/6\int\limits_{}^{}sinudu\] which is \[-\cos(u)/6 + C\] sub back in for u = 3x^2 \[-1/6\cos(3x^2)+C\]